OpenStudy (anonymous):

How would i solve: x^2-8x=9

OpenStudy (anonymous):

would i subtract 9 from both sides first?

OpenStudy (anonymous):

Yes. And set it to equal 0. x^2 - 8x - 9 = 0. Then factor.

OpenStudy (anonymous):

two numbers that multiply together to equal -9 and add up to -8 right?

OpenStudy (anonymous):

Yes. So what factors make up 9? 1, 9, and 3. Since 8 is a negative number. The numbers that multiply to 9 has to be one positive and one negative.

OpenStudy (anonymous):

3 won't work in this situation because they don't add up to 8. So try -1, 1, 9, or -9. And plug it in into the binomials: (x )(x )

OpenStudy (anonymous):

(x-1)(x+9)

OpenStudy (anonymous):

Correct!

OpenStudy (anonymous):

Now that you have it all factored: (x - 1)(x + 9) = 0 If it's asking to you solve for the zeros, or check if it's extraneous then you'd set both to 0 like: x - 1 = 0 and x + 9 = 0. And solve.

OpenStudy (anonymous):

It just says to solve the equation.

OpenStudy (anonymous):

Ok, so for the next one. x^2+2x=4 i subtract 4 from both sides. but im stuck after that.

OpenStudy (anonymous):

So then you'd rewrite it to make it simple for you: x^2 + 2x - 4 = 0.

OpenStudy (anonymous):

You can't factor out now because the only numbers to get -4 is (-2), and (2) but it does not equal to 2 when you add them. So you're going to have to use the quadratic formula for this problem and its: